The U.S. Department of Agriculture (USDA) is reviewing a proposal to pay farmers a price support of $3.00 per bushel for a given commodity. The market demand for the commodity is given by the equation QD = 16 – 4P, where Q is in billions of bushels. The market supply of the commodity is given by the equation QS = −2 + 2P. Rather than buying surplus production, the USDA will require that farmers only produce the quantity demanded at the support price. In other words, the USDA will pay farmers not to produce any surplus production at the support price. Assuming that the proposal is enacted, the total cost to the taxpayer is estimated to be __________ billion:

Multiple Choice

  • $0.00

  • $12.75

  • $9.75

  • $10.50

  • $12.00
Answer #1

Answer:

The total cost to the taxpayer is estimated to be $12 billion because USDA will buy only 4 billions of bushel of the commodity at the price of $3 at which the market is in equilibrium i.e. QD=QS=>16-4P=-2+2P=>P=3 and Q=4.

And USDA will not buy any surplus production which means USDA will buy only the quantity based on the market demand which is Q=4 billions of bushel and pay farmers at the market price which is P=$3. Thus total purchase of USDA=P*Q=$3*4=$12.

Hence, the total cost to the taxpayer is $12 billion.